OpenStudy (anonymous):

Integrate using trigonometric substitution: x ------- dx? (x^2 -4)

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

lat x=2sectheta \[\int\limits_{}^{}\frac{\sec \theta }{\tan \theta}\sec \theta \tan \theta d \theta\]

OpenStudy (anonymous):

\[\int\limits_{}^{}\sec ^{2}\theta d \theta =\tan \theta +c \]

OpenStudy (anonymous):

so the integral is what?

OpenStudy (anonymous):

\[\int\limits_{}^{}(\sec \theta)\sec \theta \tan \theta d \theta /\tan^2 \theta =\int\limits_{}^{}\sec^2 \theta d \theta /\tan \theta\]

OpenStudy (anonymous):

hold on is this just \[\int \frac{x}{x^2-4}dx\]

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

and if so why is this a trig sub?

OpenStudy (anonymous):

is it log(x^2-4)/2??

OpenStudy (anonymous):

looks like a fairly simple partial fraction problem. i think you get \[\int\frac{1}{x-2}dx + \int\frac{1}{x+2}dx\]

OpenStudy (anonymous):

oops no sorry there should be a 1/2 out front

OpenStudy (anonymous):

one half of that whole thing.

OpenStudy (anonymous):

Because that was the question I had in the list of practice questions

OpenStudy (anonymous):

oh ok. well after you make the trig sub you are going to have to switch back

OpenStudy (anonymous):

wait wait you do not even need partial fractions. just a simple u-sub

OpenStudy (anonymous):

\[u=x^2-4\] \[du=2xdx\] and you aer done in one step

OpenStudy (anonymous):

answer is \[\frac{1}{2}\ln(x^2-4)\]

OpenStudy (anonymous):

maybe the teacher meant u - sub instead of trig sub

OpenStudy (anonymous):

shine had it earlier, sorry

OpenStudy (anonymous):

Does seem easier, thanks for the help :D

OpenStudy (anonymous):

sorry about the false start. partial fractions indeed!